# A Digital twin for Diesel Engines: Operator-infused PINNs with Transfer Learning for Engine Health Monitoring

### v1

@@ 5.1 Training of multi-head neural networks
5.2 Parameter and state estimation with clean data: Few shot method
After training,, the next step is to perform PINN to identify the engine parameters (Phase-II). The output layer ($\bm{W}_{out}$, $b_{out}$) of the multi-head networks is replaced with one output and trained for a new control vector ($\bm{u}=\{u_{\delta},u_{vgt},u_{egr}\}$) and engine speed ($n_{e}$). The weights of the output layer are initialized from a random Gaussian distribution with mean zero and standard deviation $\sigma=\sqrt{2/(\text{dim}_{\text{in}}+\text{dim}_{\text{out}})}$. The weights are sorted after initialization, and it shows better and faster convergence. The biases are initialized with zero. Similar to the PINN and transfer learning method, each of the self-adaptive weights for the data loss term of $\bm{\lambda}_{p_{em}}$, $\bm{\lambda}_{p_{im}}$, $\bm{\lambda}_{\omega_{t}}$ and $\bm{\lambda}_{W_{egr}}$ are initilized from a Gaussian random variable with mean $\dfrac{\alpha X_{data}}{\max(X_{data})}$, $X=\{p_{em},p_{im},\omega_{t},W_{egr}\}$ and standard deviation $\sigma=0.1$, where $\alpha$ is a initial weight given ($10^{3}$, $10^{3}$, $10^{3}$ and $2\times 10^{3}$ for $\bm{\lambda}_{p_{em}}$, $\bm{\lambda}_{p_{im}}$, $\bm{\lambda}_{\omega_{t}}$ and $\bm{\lambda}_{W_{egr}}$ respectively). The self-adaptive weight $\lambda_{T_{em}}$ and $\lambda_{T_{1}}$ are initialized from a Gaussian random variable with mean $20$ and standard deviation 0.1. We consider a softplus mask function to the self-adaptive weights so that the weights are always positive. The unknown parameters are initialized from a Gaussian random variable with standard deviation $\sigma=0.2$ and mean as $0.7$, $2.25$, $1.4$ and $2.5$ for $\eta_{sc}$, $h_{tot}$, $A_{egrmax}$ and $A_{vgtmax}$ respectively. Furthermore, as these quantities are always positive, we consider a mask function and scaling as shown in Table LABEL:Table:Mask_and_scaling_Unknown_parameter. Also note that the trainable parameters of the problem are the weights and bias of the last layer of the neural networks, the unknown parameters of the engine and the self-adaptive weights. These are optimized using eight different Adam optimizers: one for the weights and the biases of the neural networks, one for the unknown parameters, and one for each of the six self-adaptive weights. The loss function is given by the equation 20 as discussed in section 2. The learning rate scheduler is shown in Table LABEL:Table:Learniong_rate_schedule_few_shot (in Appendix C). We would also like to mention that we only optimized the self-adaptive weights up to $20\times 10^{3}$ epoch. After that, we fixed the self-adaptive weights with the value at $20\times 10^{3}$ epochs. The size of the self-adaptive weights for $\bm{\lambda}_{p_{em}}$, $\bm{\lambda}_{p_{im}}$, $\bm{\lambda}_{\omega_{t}}$ and $\bm{\lambda}_{W_{egr}}$ are same as that of the number of data point considered for each of them. Furthermore, as discussed earlier, in order to reduce the computational cost, we perform the derivative of the output with respect to the input outside the training loop as follows,
$y=S_{p}\left(\bm{H}(t)\bm{W}_{out}+b_{out}\right)$ (27a)
$\dfrac{dy}{dt}=\dfrac{1}{1+\exp[\bm{H}(t)\bm{W}_{out}+b_{out}]}\exp\left[\bm{H}(t)\bm{W}_{out}+b_{out}\right]\dfrac{d\bm{H}(t)}{dt}\bm{W}_{out}$ (27b)
where $S_{p}$ is softplus function (refer Table 8), $\bm{H}(t)$ is the output of the last hidden layer, $\bm{W}_{out}$ and $b_{out}$ are the weights and bias of the output layer, $\dfrac{d\bm{H}(t)}{dt}$ is the derivative of the output of the last hidden layer with respect to input $t$. From Eq. 27b, the derivative term is equal to $\dfrac{d\bm{H}(t)}{dt}$, and as the hidden layers are not trained, this remains fixed and can be evaluated analytically outside the network training loop. The size of $\dfrac{d\bm{H}(t)}{dt}$ is a matrix with size $n_{t}\times n_{h}$, where $n_{t}$ is the number of input points in time, and $n_{h}$ is the number of neurons in the last hidden layer.
We study two different scenarios with clean data without additional noise in the field data (simulated) for the inverse problem and a case with additional noise. We study five minutes of engine operation and sub-divide into five independent problems, each studied independently. For each one minute, we consider 301 residual points with equal $dt=0.2$ sec. For the first case of clean data, we consider $151$ data points ($dt=0.4$ sec.) for the data loss of $p_{em}$, $p_{em}$, $\omega_{t}$, $W_{egr}$ and $T_{em}$. As discussed earlier, the size of self-adaptive weight in this case is $151\times 1$ for $\lambda_{p_{em}}$, $\lambda_{p_{im}}$, $\lambda_{\omega_{t}}$ and $\lambda_{W_{egr}}$, the size self-adaptive weight for $\lambda_{T_{em}}$ and $\lambda_{T_{1}}$ is $1\times 1$. For clean data, we consider $301$ data points ($dt=0.2$ sec) for calculating data loss for $p_{em}$, $p_{em}$, $\omega_{t}$, $W_{egr}$ and $T_{em}$. The size of the self-adaptive weights is $301$ for $\lambda_{p_{em}}$, $\lambda_{p_{im}}$, $\lambda_{\omega_{t}}$ and $\lambda_{W_{egr}}$ while the size of self-adaptive weight for $\lambda_{T_{em}}$ and $\lambda_{T_{1}}$ is $1\times 1$. We note that despite using clean data, uncertainty is included in our model due to use of dropouts in pre-trained networks. Furthermore, in order to check the effect of initialization on the accuracy, we consider $30$ independent runs, for each one minute data segment.
Table 9 shows the mean and standard deviation of the predicted unknowns for each time segment. The corresponding violin plots are shown in Figs. 12 and B.3.1 for clean data with 151 and 301 points points respectively. One of the realizations of the dynamics of the predicted gas flow states and $x_{r}$ and $T_{1}$ with 151 data points is shown in Fig. 13. Similarly, we show the corresponding intermediate state variables in Fig. 14. All these results are normalized to a scale 0-1 using Eq. 26 for confidentiality. From Table 9 and Figs. 12 and B.3.1, it can be observed that predicted parameters show a good accuracy except during the 2-3 minutes duration. As discussed earlier, this may be due to an engine idle case during this time. We also observe outliers in $h_{tot}$ prediction during 1-2 minute when 151 data points are considered, while this is reduced when we considered 301 data points.
Fig. 12: Violin plot, few-shot learning, clean data, 151 data points: of the predicted value of the unknown parameters. 151 data points are considered for each of the known parameters (for each one minute segment) and no noise is considered in this case. The violin plots are clipped at the maximum and minimum values of the predicted parameters. The results are shown normalizing the true value of the unknown parameters to be one. The dots over the violin plots show the swarm plot indicates the predicted value for individual runs. The grey region shown between the dotted lines indicates the $\pm$ 7.5% error band. The middle dotted line indicates the true value of the unknown parameter, the top dotted line indicates 1.075 times the true value while the bottom dotted line indicates 0.925 times the true value.
@@ 5.2 Parameter and state estimation with clean data: Few shot method
Table 9: Predicted unknown parameters (Few-shot learning): The mean and standard deviation of the predicted unknown parameters when predicted using Few-shot transfer learning with clean data and Gaussian noisy data. The first set of results are predicted mean $\pm$ standard deviation when 151 data points for the known variables (for each 60 sec segment) with clean data, the second set of results are when 301 data points for the known variables (for each 60 sec). The third set of results for noisy data with 301 data points for each known variables (for each 60 sec segment).
Duration (sec) $A_{egrmax}\times 10^{-4}$ $A_{vgtmax}\times 10^{-4}$ $\eta_{sc}$ $h_{tot}$ Remarks
Parameter estimates, Clean data
$0\sim 60$ $1.0073\pm 0.0046$ $1.0201\pm 0.0008$ $0.9929\pm 0.0086$ $0.9592\pm 0.0296$ Clean data, 151 data points for each known variables (for each 60 sec.).
$60\sim 120$ $0.9531\pm 0.0026$ $1.0115\pm 0.0007$ $0.9980\pm 0.0162$ $1.0559\pm 0.1541$
$120\sim 180$ $0.9512\pm 0.0238$ $0.9925\pm 0.0022$ $0.9281\pm 0.0357$ $0.7854\pm 0.1144$
$180\sim 240$ $1.0057\pm 0.0013$ $1.0186\pm 0.0006$ $0.9755\pm 0.0091$ $0.9589\pm 0.0173$
$240\sim 300$ $0.9238\pm 0.0032$ $0.9993\pm 0.0018$ $0.9965\pm 0.0043$ $0.9951\pm 0.0163$
$0\sim 60$ $1.0038\pm 0.0044$ $1.0198\pm 0.0007$ $0.9922\pm 0.0056$ $0.9599\pm 0.0207$ Clean data, 301 data points for each known variables (for each 60 sec.).
$60\sim 120$ $1.0312\pm 0.0042$ $1.0043\pm 0.0008$ $0.9958\pm 0.0111$ $1.0084\pm 0.0883$
$120\sim 180$ $0.9551\pm 0.0263$ $0.9922\pm 0.0014$ $0.9385\pm 0.0315$ $0.8205\pm 0.1001$
$180\sim 240$ $1.0057\pm 0.0013$ $1.0190\pm 0.0006$ $0.9688\pm 0.0098$ $0.9532\pm 0.0162$
$240\sim 300$ $0.9385\pm 0.0041$ $1.0024\pm 0.0006$ $0.9893\pm 0.0047$ $0.9646\pm 0.0156$
Predicted unknown, Noisy data.
@@
$0\sim 60$ $0.9699\pm 0.0653$ $1.0190\pm 0.0056$ $0.9913\pm 0.0094$ $0.9568\pm 0.0379$ Noisy data, 301 data points for each known variables (for each 60 sec.).
$60\sim 120$ $1.0255\pm 0.0748$ $0.9992\pm 0.0092$ $0.9888\pm 0.0230$ $0.9445\pm 0.1749$
$120\sim 180$ $0.5800\pm 0.0568$ $0.9520\pm 0.0165$ $0.9524\pm 0.0429$ $0.9087\pm 0.2365$
$180\sim 240$ $0.9972\pm 0.0246$ $1.0139\pm 0.0081$ $0.9554\pm 0.0276$ $0.9114\pm 0.0721$
$240\sim 300$ $0.9525\pm 0.0998$ $0.9942\pm 0.0093$ $0.9849\pm 0.0122$ $0.9519\pm 0.0390$
True value $1.000$ $1.000$ $1.000$ $1.000$
From Fig. 13, it can be observed that neural network predictions show a good agreement with the true value except $x_{r}$ and $T_{1}$. Similar to the multi-stage transfer learning cases, we observed that the effect of $x_{r}$ and $T_{1}$ on the derived variable are minimum, as can be seen from the results of $T_{e}$ and $T_{em}$ in Fig. 14. We have shown the different intermediate states by using the relationships between the primary state variables (NN output) and predicted known and constants in Fig. 14. These variables also show good agreement with the true value. As discussed earlier, we have run the model 30 times for each time zone, and the ensemble data are shown in Appendix B.3. The spikes visible in the predicted results are due to the dropout introduced in the pre-trained networks.
5.3 Parameter and state estimation with noisy data: Few shot method
In the previous section, we discussed the results for the predicted parameters and the state variable with clean data for the field variables. i.e. we did not consider any added noise to the simulated data. However, field data may be noisy in practice due to the inherent instrument noise or other reasons. Thus, to simulate a more practical scenario and the applicability of the proposed method, we consider added noise and perform PINN. We added a Gaussian noise of 3%, 3%, 1%, 5% and 3% added to the simulated data of $p_{em}$, $p_{im}$, $\omega_{t}$, $W_{egr}$ and $T_{em}$ respectively. In this case, we consider 301 data points. The other parameters are kept the same as discussed in the previous section for clean data. Before discussing further, we would like to mention that in this case, there are three different types of uncertainties: (i) uncertainty associated with the initialization of the last layer of the neural network and other parameters (e.g. unknown parameters, self-adaptive weights), (ii) uncertainty in the pre-trained network due to the dropout considered, (iii) uncertainty due to the added noise in the field variables. The mean $\pm$ standard deviation of the predicted unknown parameters for each time segment is shown in Table 9 and the corresponding violin plots are shown in Fig. 15. We observed that the standard deviation of the predicted unknown has increased compared to the clean data. This is expected due to aleatory uncertainty due to additional noise in data. Similar to the clean data case, the results in the 2$\sim$3 minute duration show a higher deviation due to engine idling scenario.
Fig. 13: Predicted states, Few-shot learning, Clean data, 151 data points: One of the predicted realization of the state variables and $x_{r}$ and $T_{1}$, for the case where clean data and 151 data points are considered for each known variables (for each one minute duration). Each 1 minute is trained independently and combined to get the total 5 minutes results. The vertical dotted blue lines separate each minute result. The results are normalized using Eq. 26 with maximum and minimum true value calculated over the 5 minute.
Fig. 14: Other variables, Few-shot learning, Clean data, 151 data points: One of the predicted realization of different variables of the diesel engine for the case where clean data and 151 data points are considered for each known variable (for each one minute duration). The variable $A_{egr}$, $T_{e}$, $T_{em}$ and $W_{t}$ are the variable which are directly depends on the unknown parameters $A_{egrmax}$, $\eta_{sc}$, $h_{tot}$ and $A_{vgtmax}$. The variable $T_{e}$ also depends on the variable $T_{1}$ and $x_{r}$.
Fig. 15: Violin plot, few-shot learning, noisy data, 301 data points: of the predicted value of the unknown parameters. 151 data points are considered for each of the known parameters (for each 1 minute) and no noise is considered. The violin plots are cut at the maximum and minimum values of the predicted parameters. The results are shown normalizing the true value of the unknown parameters to be one. The dots over the violin plots show the swarm plot indicating each predicted value. The grey colour shedding shown between the dotted lines indicates the $\pm$ 7.5% error band. The middle dotted line indicates the true value of the unknown parameter, the top dotted line indicates 1.075 times the true value while the bottom dotted line indicates 0.925 times the true value.
@@ 5.3 Parameter and state estimation with noisy data: Few shot method
Fig. 16: Predicted states, Few-shot learning, Noisy data, 301 data points. One of the predicted realization of the state variables and $x_{r}$ and $T_{1}$, for the case where clean data and 151 data points are considered for each known variables (for each one minute duration). Each 5 minute are trained independently and combined to get the total 5 minute results. The vertical dotted blue lines separate each results. The results are normalized using Eq. 26 with maximum and minimum true value calculated over the 5 minute.
We study the dynamics of the state variables when we add noise to the field variables. As discussed earlier, we would also like to mention again that we did consider any noise in the input data (i.e. input control vector $\bm{u}=\{u_{egr},u_{vgt},u_{delta}\}$ and engine speed $n_{e}$) as the input data are collected from actual engines running condition. The dynamics of one of the realizations of the predicted state variables are shown in Fig. 16. It can be observed that even with noisy data, the predicted state variables show a good agreement with the true value. We also study the intermediate state variables, and the corresponding intermediate variables are shown in Fig. 17. It is observed agreement in the predicted results except in $A_{egr}$ in 2$\sim$3 time zone. However, the dynamic pattern matches even in this time zone. This is because in the time zone, the predicted $A_{egrmax}$ deviates from the true value for this particular realization. The little noise observed in the predicted results is excepted as we are considering noisy data. The predicted empirical formulae using the pre-trained neural networks are show in Fig. B.4.3. Similar to the clean data case, we observed noise in the predicted empirical formulae due to dropout. However, in this case, the noise is more as we have considered added noise in the field data. We have shown the ensemble data in the Appendix.
@@
Fig. 17: Other variables, Few-shot learning, Noisy data, 301 data points. One of the predicted realization of different variables of the diesel engine for the case where clean data and 151 data points are considered for each known variable (for each one minute duration). The variable $A_{egr}$, $T_{e}$, $T_{em}$ and $W_{t}$ are the variable which are directly depends on the unknown parameters $A_{egrmax}$, $\eta_{sc}$, $h_{tot}$ and $A_{vgtmax}$. The variable $T_{e}$ also depends on the variable $T_{1}$ and $x_{r}$.
@@ 6 Comparison between Multi-stage transfer learning and Few-shot transfer learning
We have presented two approaches for efficiently identifying unknown parameters of a diesel engine using a hybrid PINNs based approach. In this section, we compare the two methods based on empirical evidence from our numerical experiments. We note that both methods yield comparable results in terms of the mean values of unknown parameters. However, the multi-stage transfer learning approach shows a higher variation in parameter estimation in general as compared to Few-shot transfer learning. This is attributed to the fact that the few shot transfer learning network was trained on a much larger dataset that captures a wide variety of engine operating conditions. Consequently, generalization is better with the few-shot approach as compared to the multi-stage transfer learning approach. For noisy data, we observe a much wider spread in the parameter estimation from the multi-stage transfer learning approach as compared to few-shot learning. This could be indicative of the need to retrain more layers of the transfer learning network instead of the last output layer to improve system dynamics approximation. For a particular time segment of $2-3$ minute, we observe that the estimation for $A_{egrmax}$ from the few-shot learning paradigm is less accurate as compared to the transfer learning approach. This might be because the flow through the EGR ($W_{egr}$) is very small (negligible) compared to the other time zone. Additionally, we do not observe a significant time difference in training the network with 151 or 301 data points. Any variations in computation time are transients due to network latency and the existing state of the computing node. Table 10 shows the computational time comparison between the baseline PINN model (hybrid PINN+DeepONet), multi-stage transfer learning, and few-shot learning paradigms. We note a significant improvement in computational time for the hybrid PINN methods we presented earlier compared to the standard/baseline PINN method presented in [20]. Fig. 18 shows the loss convergence for the ODE and data loss terms for different engine states and gas flows. We note that all loss terms for the two transfer learning paradigms converge earlier than the baseline PINN model to fully converge. Additionally, Fig. 19 shows the convergence plot for the four unknown engine parameters generated through an ensemble across thirty independent runs for time segment $t=3-4$ min. All unknown parameters for the transfer learning methods converge at $\approx 20K$ epochs, while for the baseline PINN model, this takes $\approx 50K$ epochs. Also, compared to the multi-stage transfer learning approach, the few-shot transfer learning method uses significantly lower computational time. This is because the few-shot method eliminates the need for automatic differentiation for the output variable. However, this approach requires a larger corpus of labeled data for the baseline training of hidden layers (body) of the PINN networks, which are trained offline. Hence, trade-offs between data availability, computational and accuracy requirements are essential for choosing one method over the other.
(a) Baseline PINN
(b) Multi-stage transfer learning
(c) Few shot learning
Fig. 18: Convergence of physics loss and data loss vs epoch for different method: (a) Baseline PINN (hybrid PINN) model, (b) Multi-stage transfer learning model, and (c) Few-shot transfer learning model. The losses plotted are without the self-adaptive weights. Losses for both the transfer learning methods converges earlier than the baseline PINN. Initial condition losses are not shown.
@@
Table 10: Computational time comparison, multi-stage transfer learning vs few-shot. Mean ($\mu$) and standard deviation ($\sigma$) of computational time are calculated from 30 independent runs for each time segment. We see a significant improvement by using the multi-stage transfer learning approach when compared against the baseline PINN (hybrid PINN) model (which we train from $0-1$ min time segment). While the hybrid multi-stage transfer learning approach reduces computational time by approximately 50%, the few-shot learning paradigm further improves the computational time by an additional 25% in comparison to the baseline PINN model. Random variation in computational time results from system level variations and latency that exist in high-performance cluster systems which were used for this study. All computations were performed with 32-bit data precision on one CPU core with Tensorflow version 1.
@@
7 Summary and future work
In this study, we developed two computationally efficient and practical PINN-based methods for identifying unknown parameters and dynamics of diesel engines. We optimized the problem at multiple levels, first introducing systematic (e.g. manufacturing) variability by adding dropout in the pre-trained networks used for estimating empirical quantities of the mean value gas flow model. Secondly, we proposed a hybrid PINN and operator-based (PINNs and DeepONet) framework as a step towards meeting real time computation requirements for field deployment. To reduce the computational time further, we introduced two transfer learning approaches: i) multi-stage transfer learning based on transferring knowledge from one section of signal to the remaining sections and ii) few-shot transfer learning where one network is trained offline on a large dataset comprising various operating conditions, and thereafter, knowledge is transferred for predicting the gas flow states for test signals. Both methods are suited for identifying unknown parameters using the known gas dynamics and appropriate sensor data, and we demonstrate this by using simulated data for different operating conditions. However, the input data to generate these simulated data are considered from actual engine running conditions. We also assess the sensitivity of our methods to field noise by adding Gaussian noise to our simulated data. Both methods perform reasonably well under both noise-free and noisy conditions. However, under noisy conditions, the estimates of unknown parameters and dynamic states show higher variability as compared to noise-free operating conditions. In terms of computational cost, the multi-stage transfer learning approach requires fine tuning on each segment of signal and is more expensive as compared to fine tuning the few-shot model that shares knowledge with a model trained on a larger data corpus. However, a higher computational cost is incurred on the training of the body (hidden layers) for the few-shot model upfront which is done offline.
One source of epistemic uncertainty in our model is the use of dropout rates in pre-trained empirical networks. In this study, we use dropout across our experiments. Understanding the sensitivity of our model’s predictions to changes in these dropout values will offer further insights into the epistemic uncertainty expected from our methods. The dropout rate should be estimated from uncertainty in the actual data so that the network prediction gives similar uncertainty. Furthermore, we also note that since dropout rate depends on network size, further study is necessary to determine the appropriate dropout values with network size. Additionally, methods to reduce the variability that we see in parameter estimation under noisy conditions is a necessary step for improving the performance of our methods when deployed on field units. Minimizing the effect of aleatoric uncertainty on the model’s prediction is critical to successful implementation of such methods in the field and we continue to investigate means for this. Another future direction may include the study of a computationally effective method for dynamical stoping of training (number of training epochs) for different operating conditions.
(a) Baseline PINN (hybrid PINN)
@@ 7 Summary and future work
(b) Multi-stage transfer learning
(c) Few-shot transfer learning

### v2

@@ 4.2 Parameter and gas flow dynamics estimation
Table 7: Predicted unknown parameters, Few-shot learning: The mean and standard deviation of the predicted unknown parameters when predicted using Few-shot transfer learning with clean data and Gaussian noisy data. The first set of results are predicted mean $\pm$ standard deviation when 151 data points for the known variables (for each 60 sec segment) with clean data, the second set of results are when 301 data points for the known variables (for each 60 sec). The third set of results for noisy data with 301 data points for each known variables (for each 60 sec segment).
Duration (min) $A_{egrmax}$ $A_{vgtmax}$ $\eta_{sc}$ $h_{tot}$
Clean data, 151 data points for each known variables (for each 60 sec.).
$0\sim 1$ $1.0073\pm 0.0046$ $1.0201\pm 0.0008$ $0.9929\pm 0.0086$ $0.9592\pm 0.0296$
$1\sim 2$ $0.9531\pm 0.0026$ $1.0115\pm 0.0007$ $0.9980\pm 0.0162$ $1.0559\pm 0.1541$
$2\sim 3$ $0.9512\pm 0.0238$ $0.9925\pm 0.0022$ $0.9281\pm 0.0357$ $0.7854\pm 0.1144$
$3\sim 4$ $1.0057\pm 0.0013$ $1.0186\pm 0.0006$ $0.9755\pm 0.0091$ $0.9589\pm 0.0173$
$4\sim 5$ $0.9238\pm 0.0032$ $0.9993\pm 0.0018$ $0.9965\pm 0.0043$ $0.9951\pm 0.0163$
Clean data, 301 data points for each known variables (for each 60 sec.).
$0\sim 1$ $1.0038\pm 0.0044$ $1.0198\pm 0.0007$ $0.9922\pm 0.0056$ $0.9599\pm 0.0207$
$1\sim 2$ $1.0312\pm 0.0042$ $1.0043\pm 0.0008$ $0.9958\pm 0.0111$ $1.0084\pm 0.0883$
$2\sim 3$ $0.9551\pm 0.0263$ $0.9922\pm 0.0014$ $0.9385\pm 0.0315$ $0.8205\pm 0.1001$
$3\sim 4$ $1.0057\pm 0.0013$ $1.0190\pm 0.0006$ $0.9688\pm 0.0098$ $0.9532\pm 0.0162$
$4\sim 5$ $0.9385\pm 0.0041$ $1.0024\pm 0.0006$ $0.9893\pm 0.0047$ $0.9646\pm 0.0156$
Noisy data, 301 data points for each known variables (for each 60 sec.).
@@
$0\sim 1$ $0.9699\pm 0.0653$ $1.0190\pm 0.0056$ $0.9913\pm 0.0094$ $0.9568\pm 0.0379$
$1\sim 2$ $1.0255\pm 0.0748$ $0.9992\pm 0.0092$ $0.9888\pm 0.0230$ $0.9445\pm 0.1749$
$2\sim 3$ $0.5800\pm 0.0568$ $0.9520\pm 0.0165$ $0.9524\pm 0.0429$ $0.9087\pm 0.2365$
$3\sim 4$ $0.9972\pm 0.0246$ $1.0139\pm 0.0081$ $0.9554\pm 0.0276$ $0.9114\pm 0.0721$
$4\sim 5$ $0.9525\pm 0.0998$ $0.9942\pm 0.0093$ $0.9849\pm 0.0122$ $0.9519\pm 0.0390$
True value (normalized)
$1.000$ $1.000$ $1.000$ $1.000$
@@
Fig. 9: Predicted states, Few-shot learning, Noisy data, 301 data points. One of the predicted realization of the state variables and $x_{r}$ and $T_{1}$, for the case where noisy data and 301 data points are considered for each known variable (for each one minute duration). Each 5 minute are trained independently and combined to get the total 5 minute results. The vertical dotted blue lines separate each result. The results are normalized using Eq. (21) with maximum and minimum true values calculated over the 5 minute.
@@
Fig. 10: Intermediate variables, Few-shot learning, Noisy data, 301 data points. One of the predicted realizations of intermediate variables of the diesel engine for the case where noisy data and 301 data points are considered for each known variable (for each one-minute duration). The variables $A_{egr}$, $T_{e}$, $T_{em}$ and $W_{t}$ are the variables which are directly dependent on the unknown parameters $A_{egrmax}$, $\eta_{sc}$, $h_{tot}$ and $A_{vgtmax}$. The variable $T_{e}$ also depends on the variable $T_{1}$ and $x_{r}$.
Fig. 11: Violin plot, few-shot learning, clean data, 151 data points: of the predicted value of the unknown parameters. 151 data points are considered for each of the known parameters (for each one minute segment) and no noise is considered in this case. Similar to the previous violin plots, these results are shown in a normalized plot with the true value of the unknown parameters to be one.
Fig. 12: Violin plot of the predicted value of the unknown parameters, when predicted using few-shot TL with 301 noisy data points considered for each of the known variables (for each 1 minute). Similar to the previous violin plots, these results are shown in a normalized plot with the true value of the unknown parameters to be one.
@@ 5 Comparison between Multi-stage transfer learning and Few-shot transfer learning
We have presented two approaches for efficiently identifying unknown parameters of a diesel engine using a hybrid PINNs based approach. In this section, we compare the two methods based on empirical evidence from our numerical experiments. We note that both methods yield comparable results in terms of the mean values of unknown parameters. The multi-stage TL approach shows a higher variation in parameter estimation in general as compared to Few-shot TL. This is attributed to the fact that the few-shot TL network was trained on a much larger dataset that captures a wide variety of engine operating conditions. Consequently, generalization is better with the few-shot TL approach as compared to the multi-stage TL approach. However, statistical analyses of the parameter estimates did not yield conclusive evidence in support of few-shot approach in predicting target parameter values across all time segments. For instance, analysing the clean data for parameter $h_{tot}$, we observe a non-significant p-value at 95% confidence interval for the one-sample t-test between 240-300s which indicates enough evidence does not exist to reject the null $H_{0}:\mu=1.0$. Hence, for $h_{tot}$, one may conclude that the few-shot approach is more reliable in certain sections of the data. But the few-shot method shows departure from the target mean value of 1.0 when estimating parameter $A_{egrmax}$ in the same time segment.
@@
For noisy data, we observe a much wider spread in the parameter estimation from the multi-stage TL approach as compared to the few-shot TL. This could be indicative of the need to retrain more layers of the multi-stage TL network instead of the last output layer to improve system dynamics approximation. For a particular time segment of 232-3 minute, we observe that the estimation for $A_{egrmax}$ from the few-shot TL paradigm is less accurate as compared to the multi-stage TL approach. This might be because the flow through the EGR ($W_{egr}$) is very small (negligible) compared to the other time zone. Additionally, we do not observe a significant time difference in training the network with 151 or 301 data points. Any variations in computation time are transients due to network latency and the existing state of the computing node. Table 8 shows the computational time comparison between the baseline PINN model (hybrid PINN+DeepONet), multi-stage TL, and few-shot TL paradigms. We observe that the proposed hybrid multi-stage TL strategy achieves nearly 50% reduction in computational time while the incorporation of few-shot learning approach yields an additional 25% improvement over the baseline PINN framework. Note that the multi-stage TL strategy still requires the full PINN framework to be trained for the initial 0-60s but learning across the remaining time segments is significantly faster. The few-shot paradigm, on the other hand requires additional training of the multi-head network but during evaluation phase, the gradient calculation for physics loss estimation is simplified leading to better computational efficiency as compared to the multi-stage TL approach.
@@
Table 8: Computational time comparison, multi-stage vs few-shot TL. Mean ($\mu$) and standard deviation ($\sigma$) of computational time are calculated from 30 independent runs for each time segment. While the hybrid multi-stage TL approach reduces computational time by approximately 50%, the few-shot learning paradigm further improves the computational time by an additional 25% in comparison to the baseline PINN model. Random variation in computational time results from system level variations and latency that exist in high-performance cluster systems which were used for this study. All computations were performed with 32-bit data precision on one CPU core with Tensorflow version 1.
@@
Fig. 13 shows the loss convergence for the ODE and data loss terms for different engine states and gas flows. We note that all loss terms for the two TL paradigms converge faster than the baseline PINN model. Additionally, Fig. 14 shows the convergence plot for the four unknown engine parameters generated through an ensemble across thirty independent runs for time segment $t=3\sim 4$ min (clean data with 151 data points) for base line PINN, multi-stage and few-shot TL. All unknown parameters for the TL methods converge at $\approx 20K$ epochs, while for the baseline PINN model, this takes $\approx 50K$ epochs. Also, compared to the multi-stage TL approach, the few-shot TL method uses significantly lower computational time. This is because the few-shot method eliminates the need for automatic differentiation for the output variable. However, this approach requires a larger corpus of labelled data for the baseline training of hidden layers (body) of the PINN networks, which are trained offline. Hence, trade-offs between data availability, computational and accuracy requirements are essential for choosing one method over the other.
@@
(a) Baseline PINN Refer to caption (b) Multi-stage transfer learning Refer to caption (c) Few shot transfer learning
Fig. 13: Convergence of physics loss and data loss vs epoch for different method: (a) Baseline PINN (hybrid PINN) model, (b) Multi-stage transfer learning model, and (c) Few-shot transfer learning model when 151 clean data is considered. The losses plotted are without the self-adaptive weights. Losses for both the transfer learning methods converge earlier than the baseline PINN. Initial condition losses are not shown.
(a) Baseline PINN (hybrid PINN) Refer to caption (b) Multi-stage transfer learning Refer to caption (c) Few-shot transfer learning